Solve y – 3y - 10y = e2x + 5x + 10x by undetermined coefficients.
Answer #1

Page: Date: y -34 -10y = e2x + osx 4106 2y = 2x tesx +lox CE = → mb-3m-10 = 0 _ ma- smt2 m -10 = 0 mcm - 5) +2cm-5) =0 (m+2)